OpenStudy (anonymous):

Henry and Layla are ordering flowers for a birthday party. Large arrangements are $75 and small arrangements are $50. They can afford to spend no more than $700, but would like to purchase more than 10 arrangements. Which is not a combination of large and small arrangements that Henry and Layla could purchase? 3 large and 8 small arrangements 4 large and 7 small arrangements 5 large and 6 small arrangements 7 large and 4 small arrangements

OpenStudy (agent0smith):

L for large, S for small. they want more than 10 total: \[\large L + S >10\] L's cost 75, S's cost $50. They can't spend more than $700: \[\large 75L + 50S \le 700\] Now just try each of 3 large and 8 small arrangements 4 large and 7 small arrangements 5 large and 6 small arrangements 7 large and 4 small arrangements to see which ones will actually work in the two inequalities above.

OpenStudy (agent0smith):

Remember, large is L and small is S, so use the numbers given and find one pair of number that does NOT work in the inequalities.

OpenStudy (anonymous):

Still a little confused :/

OpenStudy (anonymous):

I got the first one?

OpenStudy (anonymous):

Am I right?

OpenStudy (agent0smith):

3 + 8 > 10 this is true since 11 is greater than 10. \[\large 75*3 + 50*8 \le 700\] is this true or false? if it's true, it's not your answer

OpenStudy (anonymous):

its true :)

OpenStudy (agent0smith):

Move on to the next two numbers and check those then

OpenStudy (anonymous):

ohhhh I have to find out which isn't true :/

OpenStudy (agent0smith):

Yep, check each one :)

OpenStudy (anonymous):

Got it, after trying all of them, its the last one..?

OpenStudy (agent0smith):

7 large and 4 small add up to more than 11 \[\large 75*7 + 50*4 \le 700\] 75*3 + 50*8 = 725 which is NOT less than 700

OpenStudy (anonymous):

yayy

OpenStudy (agent0smith):

:)